# DLCR: A Generative Data Expansion Framework via Diffusion for Clothes-Changing Person Re-ID

### v1

@@ 3.2 Stage 2: Re-ID Training and Refinement
Our generated training set $G_{train}$ can be used to enrich the original training data $D_{train}$, improving the discriminative capabilities of any Re-ID model. One simple way to utilize $G_{train}$ is to concatenate it with ${D_{train}}$ into one dataset, then use this merged dataset to train a Re-ID model. However, the model struggles to effectively utilize $G_{train}$ when all the clothes-changed variations are introduced at once during training (discussed in supplementary). Moreover, the training time of the CC-ReID models would be heavily impacted with the generated training data being $10$x larger in size. In the supplementary, we provide a t-SNE visualization of real vs. generated images showing that the generated data exhibits a higher variance. To gently and more efficiently accommodate CC-ReID models to the diverse distribution of the generated data, we propose a progressive learning approach that sophisticatedly increases the clothing diversity of a subject as training progresses. Furthermore, the novel ability to controllably inpaint realistic clothes using our method can be extended even further than just during training. Specifically, we propose a prediction refinement strategy that uses our diffusion-based inpainting method as a unique test-time augmentation to provide a more holistic and ensemble-based decision during evaluation.
@@ 3.2.1 Progressive Learning Strategy
To mitigate impacts to training time, we gradually introduce $G_{train}$ by injecting our generated data into each training batch in a pairwise fashion. During batch construction, we pair every original training image in the batch with one of its corresponding generated samples, effectively doubling the original batch size. For brevity, consider one image from the original training set $x\in D_{train}$ and its corresponding generated samples $P^{x}=\{\hat{x}_{i}\}_{i=1}^{K}$. A generated image from $P^{x}$ can be chosen as a pair for $x$ through random sampling: $\hat{x}\sim\mathcal{U}(P^{x})$. Due to the large increase in clothing diversity and number of images in $G_{train}$, we show that incrementally introducing the clothes-changed generated images from $P^{x}$ as training progresses yields better results. Moreover, since we generate roughly $10$x more training data for each CC-ReID dataset, we found that our progressive learning strategy enables models to effectively utilize our generated data without major impacts to training time (see supplementary). Specifically, we perform progressive learning by first dividing $P^{x}$ into four distinct, equal partitions $P^{x}=P_{1}+P_{2}+P_{3}+P_{4}$. Then, during the early stages of training, only the generated images $\hat{x}\in P_{1}$ are used for sampling when finding a pair for $x$, where $\hat{x}\sim\mathcal{U}(P_{1})$. As training continues, we periodically increase the sampling range for the generated images by incorporating more partitions. For instance, $x$ is paired with a randomly sampled generated image $\hat{x}\sim\mathcal{U}(\{P_{1}+P_{2}\})$ after $N$ epochs of training have been completed. After $2N$ completed training epochs, the generated pair image is sampled by $\hat{x}\sim\mathcal{U}(\{P_{1}+P_{2}+P_{3}\})$ until eventually, every generated image in $P^{x}$ is used for sampling during batch construction after $3N$ training epochs: $\hat{x}\sim\mathcal{U}(P^{x})$. In this manner, we increase the difficulty of learning clothes-invariant features by controlling the number of clothes-changed generated images seen during the early stages of training. Therefore, the model gradually learns to ignore more clothing items as training progresses due to the increasing diversity of subject attire.
@@ 3.2.2 Prediction Refinement Strategy
While we have only discussed using DLCR for generating CC-ReID training data, we exhibit in this section that we can also leverage DLCR at test-time by employing our inpainting method (Sec. 2) to augment a given query image by changing its clothes. For a given query, we create $l$ augmented query variants. For each query variant, we then retrieve the top-$m$ similarity scores from the gallery and ensemble them to mitigate clothing bias when making a prediction. Consequently, our ensembling method yields a refined similarity score estimate of a subject ID, resulting in more accurate top-1 retrievals. Algorithm 1 describes the detailed steps of our ensembling process. Our algorithm takes a query image, the $l$ augmented versions of the query, the gallery, and the CC-ReID model as input. The first step computes the features of the query ($f^{q}$), augmented query variants $\{\hat{f}^{q}_{i}\}_{i=1}^{l}$ and gallery images $\{\hat{f}^{g}_{i}\}_{i=1}^{n}$ (Line $1$). In step two, for the original query, we retrieve the top-$m$ most similar examples, $I^{m}$, from the gallery (Lines $2$-$3$). In the fourth step, each unique subject $\{s_{i}\in S\}_{i=1}^{m}$ in $I^{m}$ is assigned an initial score. We repeat this retrieval process for each augmented query variant $\{\hat{f}^{q}_{k}\}_{k=1}^{l}$, resulting in their own top-$m$ retrievals from the gallery, $\{I_{k}^{m}\}_{k=1}^{l}$ (Line $5$). For each $I_{k}^{m}$, we update the similarity scores only for the subjects present in $S$ (Line $6$). At the end, we obtain the set of subject IDs $S$, with their corresponding refined similarity scores. Hence, for a given query image, we leverage our diffusion-based inpainting and ensembling algorithm to yield a better clothes-invariant prediction. Algorithm 1 is also extremely computationally efficient, with the entire process having a throughput of $\sim 640$ images/sec with a memory constraint of $<2$ GB.
@@ 4.1 Comparison with State-of-the-Art Methods
Table 2: Comparisons with previous CC-ReID models on the PRCC, LTCC, CCVID, and VC-Clothes datasets in the clothes-changed (CC) setting. DLCR outperforms basically every model across four datasets, with the green text denoting improvements over the CAL baseline. $\dagger$ denotes reproduced results using provided open-source code.
@@ 4.3 Performance on Out-of-Distribution Data
We show in this section that CAL+DLCR improves performance on out-of-distribution testing data as well. As previously mentioned, CAL introduces architectural changes and loss formulations, such as a clothes-classifier and clothes adversarial loss, to enforce clothes-agnostic learning. However, discriminative changes like these may only enable a model to perform better on data that it was trained on, while poorly generalizing when evaluated on unseen data. To this point, we show results in Table 4 where we take a pre-trained CAL model and a pre-trained CAL+DLCR model on the LTCC dataset, and evaluate them in a zero-shot fashion on other CC-ReID datasets (PRCC, LaST [45], and VC-Clothes). Top-1 accuracy gains ranging from $4\%$ to $8\%$ are seen across all three CC-ReID datasets, with similar gains on mAP, ranging from $\approx 1\%$ to $6\%$. Therefore, DLCR both enhances a model’s performance on a single dataset and even on out-of-distribution testing data.
Table 4: Comparisons between CAL+DLCR and CAL models pre-trained on LTCC when evaluated on other datasets. CAL+DLCR consistently beats CAL when evaluated on other datasets in a zero-shot fashion, exhibiting DLCR’s discriminative benefits extend beyond just the training dataset.
@@ 5 Conclusion
In this work, we proposed the first use of diffusion for generative data expansion in CC-ReID. Through diffusion inpainting, DLCR is capable of controllably generating completely new clothing items onto a subject while retaining person-specific information. Guided by clothing prompts extracted using foundational language models, DLCR synthesizes additional training data to increase the clothing diversity in a CC-ReID dataset. During CC-ReID training, a progressive learning strategy is applied to efficiently utilize this generated data. To better refine a model’s prediction during inference, we employ a test-time augmentation. We demonstrate that DLCR significantly improves CC-ReID performance and achieves SOTA results across multiple benchmark datasets. Finally, we release all of our generated data for the benchmark datasets used in this paper for future research.

### v2

@@ 3.2 Stage 2: Re-ID Training and Refinement
Our generated training set $G_{train}$ can be used to enrich the original training data $D_{train}$, improving the discriminative capabilities of any Re-ID model. One simple way to utilize $G_{train}$ is to concatenate it with ${D_{train}}$ into one dataset, then use this merged dataset to train a Re-ID model. However, the model struggles to effectively utilize $G_{train}$ when all the clothes-changed variations are introduced at once during training (discussed in supplementary). Moreover, the training time of the CC-ReID models would be heavily impacted with the generated training data being 10×10$\times$10 × larger in size. In the supplementary, we provide a t-SNE visualization of real vs. generated images showing that the generated data exhibits a higher variance. To gently and more efficiently accommodate CC-ReID models to the diverse distribution of the generated data, we propose a progressive learning approach that sophisticatedly increases the clothing diversity of a subject as training progresses. Furthermore, the novel ability to controllably inpaint realistic clothes using our method can be extended even further than just during training. Specifically, we propose a prediction refinement strategy that uses our diffusion-based inpainting method as a unique test-time augmentation to provide a more holistic and ensemble-based decision during evaluation.
@@ 3.2.1 Progressive Learning Strategy
To mitigate impacts to training time, we gradually introduce $G_{train}$ by injecting our generated data into each training batch in a pairwise fashion. During batch construction, we pair every original training image in the batch with one of its corresponding generated samples, effectively doubling the original batch size. For brevity, consider one image from the original training set $x\in D_{train}$ and its corresponding generated samples $P^{x}=\{\hat{x}_{i}\}_{i=1}^{K}$. A generated image from $P^{x}$ can be chosen as a pair for $x$ through random sampling: $\hat{x}\sim\mathcal{U}(P^{x})$. Due to the large increase in clothing diversity and number of images in $G_{train}$, we show that incrementally introducing the clothes-changed generated images from $P^{x}$ as training progresses yields better results. Moreover, since we generate roughly 10×10$\times$10 × more training data for each CC-ReID dataset, we found that our progressive learning strategy enables models to effectively utilize our generated data without major impacts on training time (see supplementary). Specifically, we perform progressive learning by first dividing $P^{x}$ into four distinct, equal partitions $P^{x}=P_{1}+P_{2}+P_{3}+P_{4}$. Then, during the early stages of training, only the generated images $\hat{x}\in P_{1}$ are used for sampling when finding a pair for $x$, where $\hat{x}\sim\mathcal{U}(P_{1})$. As training continues, we periodically increase the sampling range for the generated images by incorporating more partitions. For instance, $x$ is paired with a randomly sampled generated image $\hat{x}\sim\mathcal{U}(\{P_{1}+P_{2}\})$ after $N$ epochs of training have been completed. After $2N$ completed training epochs, the generated pair image is sampled by $\hat{x}\sim\mathcal{U}(\{P_{1}+P_{2}+P_{3}\})$ until eventually, every generated image in $P^{x}$ is used for sampling during batch construction after $3N$ training epochs: $\hat{x}\sim\mathcal{U}(P^{x})$. In this manner, we increase the difficulty of learning clothes-invariant features by controlling the number of clothes-changed generated images seen during the early stages of training. Therefore, the model gradually learns to ignore more clothing items as training progresses due to the increasing diversity of subject attire.
@@ 3.2.2 Prediction Refinement Strategy
While we have only discussed using DLCR for generating CC-ReID training data, we hereby explain that we can also leverage DLCR at test-time by employing our inpainting method (Sec. 2) to augment a given query image by changing its clothes. For a given query, we create $l$ augmented query variants. For each query variant, we then retrieve the top-$m$ similarity scores from the gallery and ensemble them to mitigate clothing bias when making a prediction. Consequently, our ensembling method yields a refined similarity score estimate of a subject ID, resulting in more accurate top-1 retrievals. Algorithm 1 describes the detailed steps of our ensembling process. Our algorithm takes a query image, the $l$ augmented versions of the query, the gallery, and the CC-ReID model as input. The first step computes the features of the query ($f^{q}$), augmented query variants $\{\hat{f}^{q}_{i}\}_{i=1}^{l}$ and gallery images $\{\hat{f}^{g}_{i}\}_{i=1}^{n}$ (Line $1$). In step two, for the original query, we retrieve the top-$m$ most similar examples, $I^{m}$, from the gallery (Lines $2$-$3$). In the fourth step, each unique subject $\{s_{i}\in S\}_{i=1}^{m}$ in $I^{m}$ is assigned an initial score. We repeat this retrieval process for each augmented query variant $\{\hat{f}^{q}_{k}\}_{k=1}^{l}$, resulting in their own top-$m$ retrievals from the gallery, $\{I_{k}^{m}\}_{k=1}^{l}$ (Line $5$). For each $I_{k}^{m}$, we update the similarity scores only for the subjects present in $S$ (Line $6$). At the end, we obtain the set of subject IDs $S$, with their corresponding refined similarity scores. Hence, for a given query image, we leverage our diffusion-based inpainting and ensembling algorithm to yield a better clothes-invariant prediction. Algorithm 1 is also extremely computationally efficient, with the entire process having a throughput of $\sim 640$ images/sec with a memory constraint of $<2$ GB.
@@ 4.1 Comparison with State-of-the-Art Methods
Table 2: Comparisons with previous CC-ReID models on the PRCC, LTCC, CCVID, and VC-Clothes datasets in the clothes-changed (CC) setting. DLCR essentially outperforms every model across four datasets, with the green text denoting improvements over the CAL baseline. $\dagger$ denotes reproduced results using provided open-source code.
@@ 4.3 Performance on Out-of-Distribution Data
We show in this section that CAL+DLCR improves performance on out-of-distribution testing data as well. As previously mentioned, CAL introduces architectural changes and loss formulations, such as a clothes-classifier and clothes adversarial loss, to enforce clothes-agnostic learning. However, discriminative changes like these may only enable a model to perform better on data that it was trained on, while poorly generalizing when evaluated on unseen data. To this point, we show results in Table 4 where we take a pretrained CAL model and a pretrained CAL+DLCR model on the LTCC dataset, and evaluate them in a zero-shot fashion on other CC-ReID datasets (PRCC, LaST [45], and VC-Clothes). Top-1 accuracy gains ranging from $4\%$ to $8\%$ are seen across all three CC-ReID datasets, with similar gains on mAP, ranging from $\approx 1\%$ to $6\%$. Therefore, DLCR both enhances a model’s performance on a single dataset and even on out-of-distribution testing data.
Table 4: Comparisons between CAL+DLCR and CAL models pretrained on LTCC when evaluated on other datasets. CAL+DLCR consistently beats CAL when evaluated on other datasets in a zero-shot fashion, exhibiting DLCR’s discriminative benefits extend beyond just the training dataset.
@@ 5 Conclusion
In this work, we proposed the first use of diffusion for generative data expansion in CC-ReID. Through diffusion inpainting, DLCR is capable of controllably generating completely new clothing items onto a subject, while retaining person-specific information. Guided by clothing prompts extracted using foundational language models, DLCR synthesizes additional training data to increase the clothing diversity in a CC-ReID dataset. During CC-ReID training, a progressive learning strategy is applied to efficiently utilize this generated data. To better refine a model’s prediction during inference, we employ a test-time augmentation. We demonstrate that DLCR significantly improves CC-ReID performance and achieves SOTA results across multiple benchmark datasets. Finally, we release all of our generated data for the benchmark datasets used in this paper for future research.